### v1

# CzechLynx: A Dataset for Individual Identification and Pose Estimation of the Eurasian Lynx

Abstract
We introduce CzechLynx, the first large-scale, open-access dataset for individual identification, 2D pose estimation, and instance segmentation of the Eurasian lynx (Lynx lynx). CzechLynx includes more than 30k camera trap images annotated with segmentation masks, identity labels, and 20-point skeletons and covers 219 unique individuals across 15 years of systematic monitoring in two geographically distinct regions: Southwest Bohemia and the Western Carpathians. To increase the data variability, we create a complementary synthetic set with more than 100k photorealistic images generated via a Unity-based pipeline and diffusion-driven text-to-texture modeling, covering diverse environments, poses, and coat-pattern variations. To allow testing generalization across spatial and temporal domains, we define three tailored evaluation protocols/splits: (i) geo-aware, (ii) time-aware open-set, and (iii) time-aware closed-set. This dataset is targeted to be instrumental in benchmarking state-of-the-art models and the development of novel methods for not just individual animal re-identification.
1 Background & Summary
Understanding species movement, population structure, and habitat connectivity is critical for effective conservation, especially for wide-ranging, protected carnivores. In Europe, the Eurasian lynx (Lynx lynx) exemplifies these challenges, being ecologically important, legally protected, and indicative of broader biodiversity concerns. This medium-sized, solitary, and territorial carnivore lives at low population densities, with substantial home ranges estimated at 282.9$\pm$42.4 km2 for males and 147.6$\pm$50.3 km2 for females in Central Europe [1]. Although primarily inhabiting forested areas and rugged terrains, the lynx demonstrates notable adaptability by using refuge areas to mitigate human pressures [2] and can disperse across Central Europe’s cultural landscapes, covering distances greater than 98 km from source populations [3]. Extensive and long-term monitoring efforts through camera-trapping have generated rich datasets, valuable not only for ecological research, such as estimating population densities and demographic parameters [4, 5, 6, 7], but also for the development and evaluation of novel machine learning methods. These datasets support progress in computer vision tasks, including fine-grained recognition, individual re-identification, and open-set recognition, enhancing the accuracy and efficiency of wildlife monitoring. This creates a mutually beneficial feedback loop: ecological data fuels machine learning research, and machine learning makes processing raw conservation data easier. Consolidating and standardizing these datasets for machine learning unlocks their full potential for both domains. Several studies have already been published based on the available data.
Straka et al. [8] used part of CzechLynx dataset to develop methods for 2D pose estimation tailored to endangered species. Their study focused on addressing the challenges of limited data availability by combining real images from our Central European lynx monitoring project with synthetically generated data. To ensure a robust cross-regional evaluation, they constructed training and testing datasets from two geographically distinct regions and two lynx populations: Southwest Bohemia (a Czech, Bavarian, and Austrian population) and the Western Carpathians (a Carpathian population). Using our dataset, they fine-tuned and benchmarked various state-of-the-art models, including HRNet-w32[9], and demonstrated that incorporating real lynx data significantly improved pose estimation accuracy. Their work highlights the importance and effectiveness of high-quality, species-specific datasets like ours in advancing computer vision techniques for wildlife conservation.
Picek et al. [10] used the CzechLynx dataset to evaluate a method for individual animal identification, which models foreground and background information separately. They focused on the challenges of re-identifying lynx individuals over time and across different locations, using images from the long-term camera-trap monitoring project in Central Europe. The dataset enabled them to demonstrate how incorporating spatial and temporal priors, along with their Per-Instance Temperature Scaling (PITS), improves identification accuracy, particularly when animals appear in new areas or exhibit changes in appearance over time.
Dula et al. [6] and Palmero et al. [7] used some parts of the CzechLynx dataset to estimate population densities by spatial capture–recapture (SCR) models and other demographic parameters of the Eurasian lynx in the West Carpathians in the Czech-Slovakia borderland and in the Bohemian Forest Ecosystem: Bavarian Forest National Park in Germany and the Šumava National Park in the Czech Republic. By identifying lynx individuals, Dula et al. [6] estimated apparent survival, transition rates, and population turnover. Fluctuating densities and high turnover rates indicated human-caused mortality, which could limit population growth and the dispersal of lynx to other adjacent areas, thus undermining the favourable conservation status of the Carpathian population.
1.1 Related Work
Existing animal-focused datasets cover a variety of species and provide foundational resources for developing and evaluating methods for individual animal identification. However, there is a geographic bias, with most publicly accessible datasets originating from North America and Africa, while European datasets remain less commonly shared due to data privacy and conservation concerns. Recent initiatives have aimed to create datasets specifically for the identification of individuals [11]. Currently available datasets include a wide array of mammals (e.g., zebras, belugas, whale sharks, and elephants), carnivores (e.g., leopards, hyenas, and tigers), and reptiles (e.g., turtles). However, datasets featuring large carnivores listed on the EU red lists are notably underrepresented. A summary of publicly available datasets and their relevant statistics is provided in Table 2, and non-public datasets are listed in Table 1. Brief descriptions of selected publicly accessible datasets are included below.
Individuals Images
Bears[12] 132 4,674
Dogs[13] 60 625
Jaguars 1[14] 16 176
Lynxes[15] 51 252
Polar bears 2[16] 15 42
Ocelots[17] N/A 503
Jaguars 2[17] N/A 680
Cheetahs[18] N/A N/A
Polar bears 1[19] N/A N/A
Table 1: Non-public animal re-id datasets.
The ATRW dataset [20] comprises 8,076 videos collected from ten Zoos in China and covers 92 Amur tigers (Panthera tigris). Each image is annotated with bounding boxes, identity labels, and pose keypoints, allowing development of re-identification and pose estimation methods for controlled environments.
The Hyena ID 2022 and Leopard ID 2022 datasets [21], developed by the Botswana Predator Conservation Trust, contain 3,129 images of 256 individual spotted hyenas (Crocuta crocuta) and 6,795 images of 430 African leopards (Panthera pardus), respectively. Images were taken in natural settings and annotated with bounding boxes and viewpoints, making these datasets valuable for testing in-the-wild generalization performance for individual animal re-id.
The LionData dataset [22] was compiled as part of the Mara Masai project in Kenya and includes 750 images of 98 lions (Panthera leo). Each image was manually annotated to support training and evaluation of re-id models under natural variability, such as lighting changes and different camera angles.
The PolarBearVidID dataset [23] covers 13 individual polar bears recorded in 6 German zoos. It contains approximately 138k images extracted from 1,431 video sequences. To reduce the risk of background overfitting caused by fixed camera positions, the individuals were cropped from the original photographs.
Name Species Records Individuals Span BBoxes Masks Pose In-the-wild
PolarBearVidID [23] Ursus maritimus 1,431   13 unknown $\bullet$ $\circ$ $\circ$ $\circ$
LionData [22] Panthera leo 750   94 unknown $\circ$ $\circ$ $\circ$ $\bullet$
Hyena ID 2022 [21] Crocuta crocuta 3,129 256 unknown $\bullet$ $\circ$ $\circ$ $\bullet$
Leopard ID 2022 [21] Panthera pardus 6,806 430 unknown $\bullet$ $\circ$ $\circ$ $\bullet$
ATRW [20] Pantheria tigris 8,076   92 1 day $\bullet$ $\bullet$ $\circ$ $\circ$
CzechLynx (our) Lynx lynx 37,440 219 15 years $\bullet$ $\bullet$ $\bullet$ $\bullet$
Table 2: All publicly available datasets with large carnivores. datasets based on video recordings. excluding synthetic data.
2 Methods
This section outlines the procedures used to create and prepare the CzechLynx dataset, including detailed descriptions of data acquisition, annotation, and preprocessing steps. The dataset combines data collected from long-term camera trapping surveys across two regions with synthetic samples generated to enhance variability. Where applicable, previously published methods are cited, and only novel technical processes are described in detail.
2.1 Dataset Collection
The CzechLynx dataset comprises data collected from two independent conservation projects operating within distinct regions of the Eurasian lynx’s (Lynx lynx) Central European range. Each project used its own field methodology and data management pipeline. To maintain transparency and reproducibility, we describe the sources separately below.
Carnivore Conservation Programme, Friends of the Earth Czech Republic (FoE CZ), and Department of Forest Ecology, Mendel University in Brno: This source includes data collected through long-term monitoring in the Western Carpathians (since 2009) and Southwest Bohemia (since 2015). The datasets are labeled "FoE CZ - The Western Carpathians" and "FoE CZ - Southwest Bohemia" in the following text. Camera traps were deployed on forest roads, trails, mountain ridges, as well as marking and resting sites, identified via snow-tracking, signs of occurrence, or telemetry research [24, 1]. From 2015 onward, the Western Carpathians area followed a systematic camera-trapping and spatial capture-recapture model to estimate lynx densities and evaluate density fluctuations, apparent survival, transition rate, and individual turnover during five consecutive seasons [6]. Earlier data and those collected by opportunistic camera-trapping allowed for estimation of minimum population size, social structure, and evidence of reproduction. Cameras used included both white-flash and infrared models (video-enabled). Most individuals were manually identified based on distinctive coat patterns visible on their flanks, forelimbs, and hindlimbs [6].
Šumava National Park: This source includes data collected from the areas of Šumava National Park and the Protected Landscape Area in southwestern Bohemia. The data were gathered through two long-term monitoring projects conducted in cooperation with the Šumava National Park Administration. The first project, started in 2009, covers roughly the northwestern two-thirds of Šumava National Park. Its main goal is to obtain yearly density estimates of independent lynx individuals living in the core area of the Bohemian-Bavarian-Austrian lynx population. Within the scope of the first project, data have been collected using a 2.7$\times$2.7 km grid. White-flash camera traps were installed seasonally in every second grid cell at suitable locations along forest paths, roads, and trails. Additional details about this methodology can be found in studies by Weingarth et al. [25] and Palmero et al. [26]. The second project (initiated in 2017) aims to provide as complete an overview as possible of the entire Bohemian-Bavarian-Austrian lynx population throughout its full range. Data for this project were collected year-round at various sites, including forest paths, roads, trails, and lynx scent-marking locations, using both white-flash and infrared video-enabled camera traps. These cameras were set up continuously based on a 10$\times$10 km grid (ETRS LAEA 5210). In each grid cell predominantly within the Šumava National Park and Protected Landscape Area, between 4 to 8 camera sites were continuously monitored throughout the year. Further information about this project is available in a separate report [27].
2.2 Raw Data Pre-processing
The pre-processing of raw camera trap data began with the manual removal of empty images, which frequently resulted from false triggers caused by wind, moving vegetation, or fluctuating light conditions. Once these non-informative images were discarded, the remaining photographs were screened by a team of experts and trained volunteers, who classified the content by species, including instances involving humans or vehicles. Only images positively identified as containing Eurasian lynxes were retained for further analysis. Selected images underwent a second round of expert review by specialists with extensive experience in long-term lynx monitoring. Using the species’ distinctive coat patterns, especially on the hind limbs, forelimbs, and flanks, experts manually identified individual lynxes. In each study area (Southwest Bohemia and the Western Carpathians), at least three well-trained local observers conducted this detailed identification process, supported by a cross-verification aligned with established camera-trapping standards [6, 28]. The final set of verified images, each linked to a specific, individually recognized lynx, provided the essential foundation for all subsequent annotation and analysis steps.
2.3 Video Pre-processing
The CzechLynx dataset includes data collected from various camera trap models, some of which capture still images while others record short video sequences during detection events. Still cameras typically produce image bursts, usually three frames per trigger, while video-capable traps generate longer recordings. Since manually annotating every frame in a video is not feasible, we implemented a method to automatically extract a small number of informative frames from each clip.
The selection process is based on automated detection using MegaDetector [29], a YOLOv5-based object detection model. Each frame of the video is analyzed, and those with a detection confidence below 0.7 are discarded. From the remaining frames, the first and last with valid detections are identified, and the interval between them is divided into three equal parts. One frame is then selected from each segment – specifically, the frame with the largest detected bounding box – ensuring that the chosen images capture the animal prominently and are distributed across the video timeline. Selected frames serve as representative samples for downstream tasks such as individual identification and pose estimation. The following criteria guided the frame selection process: (i) the animal’s full body should be visible; (ii) the animal should be captured from a side view, where identifying coat patterns are most apparent; (iii) the animal should be close enough to the camera to reveal fine details without any part being cropped; and (iv) the selected frames should be temporally spaced to capture variation in pose and movement. Figure 1 illustrates the selected frames and the relative bounding box areas across the video sequence.
Figure 1: Example of frame selection from video sequences based on bounding box area and temporal spacing. The video is divided into three segments (visualized as red lines), and one frame with the highest relative area is selected from each segment.
2.4 Dataset Annotation
All extracted video frames and images from camera traps were further processed, i.e., we manually annotated the data for three computer vision tasks: instance segmentation, individual identification, and animal pose estimation. We used Computer Vision Annotation Tool (CVAT), an open-source platform for manual labeling, whose web-based interface allows us to efficiently annotate polygons, masks, key points, and attribute tagging. The data for instance segmentation were annotated using a semi-automated approach based on the Segment Anything Model (SAM) [30]. Human annotators used positive and negative points to prompt SAM to generate initial masks, which were (if needed) manually fixed in CVAT to outline each lynx accurately.
The data targeted for animal pose estimation were also annotated using a semi-automated approach. Initial keypoint predictions were generated with a pre-trained AnimalPose model [31] and subsequently updated manually in CVAT if needed. Up to 20 key points were assigned to each individual based on the MMPose animal skeleton specification, covering joints, facial features, and the tail base. The detailed description of all keypoints and sample poses is described in Table 3 and Figure 2. In cases of partial visibility, some key points were left unannotated if they were not visible in the image. In addition to spatial labels, we also annotate image descriptive metadata, including the lynx’s viewpoint (e.g., left flank, right flank, frontal, or rear), an identifiability flag indicating whether individual recognition is possible.
Keypoint No. 1 2 3 4 5 6 7 8 9 10
[…]
Keypoint No. 11 12 13 14 15 16 17 18 19 20
[…]
Table 3: Keypoint definitions used for pose estimation.
Figure 2: Lynx lynx pose annotations. 20-keypoint skeletal structure based on AnimalPose standard.
2.5 Synthetic Data Generation
Synthetic data has proven valuable in improving performance in many scenarios, especially when real-world data is scarce, difficult to collect, or lacks sufficient diversity [32, 8, 33]. In the context of individual animal identification or pose estimation, however, generating synthetic data that accurately captures the complexity of real-world conditions remains a significant challenge. Despite its usefulness, existing methods for generating synthetic animal pose data often struggle with realism in appearance, motion, and environmental context, which can lead to reduced accuracy in downstream applications [34, 35, 36]. To complement the real data and broaden the variability in both poses and textures, we created a synthetic dataset using rendered 2D images of a 3D lynx model. Our generation pipeline, built with the Unity game engine, enables the creation of highly realistic synthetic samples that reflect multiple individuals of the Lynx lynx species across a wide range of conditions. By leveraging detailed environmental modeling, including vegetation, terrain, and lighting, along with lifelike animations and diverse camera setups, we ensure the production of high-fidelity synthetic images that resemble real-world scenarios.
1:Blank 3D mesh of animal, set of textures (each with probability of being selected $p_{t}$), set of animations (each with probability of being selected $p_{a}$)
2:procedure Setup
3:    Setup 3D environment
4:    Prepare a set of auxiliary animations
5:    Prepare a set of textures
6:procedure Simulation
7:    for each 3D environment and scenario do
8:         while scenario is running do
9:             With probability $p_{t}$ select a texture for the 3D mesh
10:             With probability $p_{a}$ trigger auxiliary animations
11:             Capture and store a 2D snapshot every $N$-th frame
Pipeline: We propose a pipeline to generate synthetic animal images with rich metadata, including precise pose, segmentation masks, and individual annotations. The process begins with the construction of a 3D environment in Unity, where we simulate various behavioral scenarios. This involves modeling the animal in 3D with high-resolution textures and animations, as well as adjusting environmental elements to introduce naturalistic variability. We include an animal moving through a predefined environment using a walking animation that follows multiple fixed points. An auxiliary sitting animation is randomly triggered approximately once per path between two points to include some pose variability. A single camera captures 2D snapshots of the scene, recording metadata such as pose keypoints. The camera placement ensures that all animal keypoints remain visible and unobstructed throughout the simulation. By integrating these components, we create a dynamic simulation that closely mimics real-world animal behavior. To generate multiple identities, each frame uses a randomly selected lynx texture, with each texture corresponding to a unique individual. The complete procedure is outlined in Algorithm 11.
Data generation pipeline: Our data generation process is designed to closely replicate real-world camera trap photography. We begin by manually selecting four points in each scene, some of which may be outside the camera’s field of view. The animal model then moves randomly between these points. As it walks, we randomly select $n$stop" points along the path. At each "stop", we modify both the scene and the model’s pose. Scene adjustments include selecting different environment variations (e.g., tree types, grass, snow) and making slight changes to the camera viewpoint (e.g., rotation). The model’s pose is altered by rotating it and transitioning between available animations. If a texture variation is provided, it is also randomly applied. We save two data samples: one before any adjustments and another after modifications. If the majority of the model is outside the camera’s view, the sample is discarded.
Animal model: We use a textured model of Lynx lynx as our animal model. With proper setup, other animal models can also be used. Unfortunately, the model initially included only a walking animation. To increase pose diversity, we manually created an additional sitting animation. The model’s skeletal structure includes 20 keypoints and follows the Animal Pose dataset. This allows for training on both real and synthetic datasets using the same model and enables direct performance comparisons. Additionally, we capture the model’s bounding box, inferred from its mesh.
Textures: For the 3D animal meshes used in the simulation, we produce an alternative set of textures. Rather than manually constructing each texture, we use diffusion-based models [37, 38, 39, 40], which have proven highly effective for various computer vision tasks, including texture generation for 3D meshes and enables the creation of textures directly based on text descriptions. In our pipeline, we employ the Paint-It model [40], a text-to-texture model that generates textures for blank 3D meshes based on textual prompts. To generate realistic and varied textures for our models, we use prompts such as lynx and bobcat, which are often used interchangeably in the context of image labels. Using these prompts allows the model to produce different textures that reflect the unique features of these animals. However, as shown in the first row of Figure 3, these basic prompts did not always result in diverse textures. To increase the variability, we introduced an additional 24 prompts – 12 for lynx and 12 for bobcat. These prompts provided detailed descriptions of common feline fur patterns, incorporating terms such as spotted, tabby, and marbled. As shown in the second row of Figure 3, the inclusion of these detailed prompts significantly increased the variability of the generated textures without causing significant qualitative distortion. Overall, we produced 300 unique textures.
# Samples Full text prompt
15+15 A lynx / bobcat.
$15+15$ A lynx / bobcat in natural habitat.
$10+10$ A Spotted lynx / bobcat with small, evenly spaced black spots scattered across its beige fur, showcasing a classic spotted pattern.
$10+10$ A Striped lynx / bobcat featuring thin, dark stripes running along its back and sides, adding a sleek and striking appearance to its coat.
$10+10$ A Marbled lynx / bobcat with a marbled pattern, characterized by swirling patches of lighter and darker fur, creating a mesmerizing and intricate design.
$10+10$ A Tabby lynx / bobcat displaying tabby markings, with bold, dark stripes running down its sides and a distinctive shape on its forehead, giving it a distinctive and recognizable look.
$10+10$ A Rosette lynx / bobcat with rosette-shaped markings, consisting of dark outlines surrounding clusters of lighter fur, resembling the spots of a leopard and adding a touch of elegance to its appearance.
$10+10$ A Calico lynx / bobcat with a calico pattern, featuring patches of different colors such as black, brown, and beige, creating a unique and eye-catching mosaic on its coat.
$10+10$ A Ticked lynx / bobcat exhibiting a ticked pattern, with individual hairs displaying alternating bands of light and dark colors, resulting in a subtly textured and speckled appearance.
$10+10$ A Brindle lynx / bobcat with a brindle pattern, characterized by irregular streaks and stripes of dark fur on a lighter background, adding depth and dimension to its coat.
$10+10$ A Banded lynx / bobcat displaying bands of color across its body, with distinct horizontal stripes or bands of darker fur breaking up its coat into segments.
$10+10$ A Patchwork lynx / bobcat with a patchwork pattern, featuring irregular patches of different colors and shapes distributed across its fur, creating a unique and eclectic appearance.
$10+10$ A Mackerel lynx / bobcat with a mackerel pattern, characterized by vertical stripes along its body, resembling the markings of a tiger, adding a sleek and dynamic quality to its coat.
$10+10$ A Tortoiseshell lynx / bobcat with a tortoiseshell pattern, featuring a mix of black, orange, and brown patches scattered across its fur in a unique and intricate design.
Lynx (seed 0)
Lynx (seed 1)
Bobcat (seed 0)
Bobcat (seed 1)
Striped lynx
Tabby lynx
Marbeled bobcat
Rosette bobcat
Figure 3: Examples of generated textures given various text prompts. The prompt in each row is used either with lynx or bobcat terms. The first column shows how many times each prompt (in both lynx and bobcat variant) was used with different seeds. In total, 300 textures were generated using 28 various text prompts.
Environment: Our goal is to generate synthetic data that closely resembles real images. To achieve this, we use highly realistic, freely available assets from Unity’s Book Of The Dead, including terrain, trees, logs, etc. We selected several real-world scenes as references and recreated four photorealistic scenes based on them. To enhance data variability, each scene is generated in multiple versions, differing in tree types, grass, and snow coverage. See comparison of real and synthetic scenes in Figure 4.
Figure 4: Selected synthetic data samples. Inspired by the real camera trap views, we have created four highly realistic scenes. All scenes are made publicly available for further use.
3 Data Records
The CzechLynx dataset includes real camera trap photographs and synthetic samples of the Eurasian lynx (Lynx lynx), organized around three computer vision tasks: individual identification, pose estimation, and instance segmentation. It spans more than 15 years and covers two regions in Central Europe (Southwest Bohemia and the Western Carpathians) and contains over 40,000 manually verified and labeled images. Detailed statistics on image counts, identity coverage, and geographic distribution are listed in Table 4. All images are in JPEG format, with metadata provided in structured JSON and CSV files. The metadata includes bounding boxes, 20-point anatomical keypoints, individual identity labels, observation timestamps, and camera trap locations, to name a few. The folder structure is task-specific, with separate directories for real and synthetic data, further divided by task (e.g., pose/, reid/, segmentation/) and corresponding annotations. To allow standardized development and evaluation for all use-cases, i.e., instance segmentation, animal re-identification and pose estimation, we provide multiple sub-datasets and appropriate training / validation / test splits. All are described in detail below. The dataset is available on Kaggle.
Metadata: Additional metadata are provided for a vast majority of images in the CzechLynx dataset and contain info about the origin of the data, temporal context (e.g., observation date, relative age since first sighting, and encounter sequence), spatial context (e.g., 10$\times$10 km grid-cell code, location and trap identifiers, and GPS coordinates), phenotypic annotation, and dataset partitioning flags (i.e., geo-aware, time-open, and time-closed splits). These annotations support flexible filtering and grouping by identity, time, space, and experimental split, thereby enabling reproducible, domain-aware evaluation of re-identification, pose estimation, and segmentation models. See Table 5 for a detailed description of all metadata fields.
Source # Images # Observations # Individuals Localities # Period
FoE CZ – The Western Carpathians 17,689 5,262 69 347 2009–2024
FoE CZ – Southwest Bohemia   6,442 3,374 82 74 2015–2023
Šumava National Park 13,309 9,296 106 322 2016–2024
Total 37,440 17,932 219 743 2009–2024
Table 4: Summary of data from individual sources. The FoE CZ – Southwest Bohemia and Šumava National Park datasets overlap, resulting in 38 shared individual identities.
Metadata Description
Source The provider of the data. The string beskydy, sumava, or nps correspond to FoE CZ - The Western Carpathians, FoE CZ - Southwest Bohemia, and Šumava National Park, respectively
Unique name The unique identification of Lynx Lynx individual. The format is lynx_<integer> .
Path The relative path to the file in the dataset.
Date Date when the animal was observed in yyyy-mm-dd format.
Relative age The age derived from the difference between the actual date and the first observation of the individual in the dataset.
Encounter ID of unique sequence of images in the same camera trap location.
Coat pattern Describes the lynx’s coat pattern where available with values marbled and spotted. Present in over one-third of the images.
Cell code This column contains 10 km × 10 km grid-cell identifiers in the ETRS89-LAEA (EPSG:3035) pan-European coordinate system. Each entry has the form 10kmE<easting_index>N<northing_index>
Location The string identifier of the location.
Trap ID Unique identification of camera trap.
Geo-aware split Train / test split. The images from one location belong exclusively to either.
Split Time-open Train / test split. The individuals unseen in the train split are included in the test split.
Split Time-closed Train / test split. All individuals in the test split are included in the train split.
Latitude, Longitude WGS84 coordinates representing the center of the 10$\times$10 km grid cell where the observation occurred.
Table 5: Metadata fields and their definitions. For almost all images in the CzechLynx dataset, we include info about provenance, identity, temporal context, spatial context, phenotypic annotation, and dataset-split assignment.
3.1 Task-specific subsets
The CzechLynx dataset is organized into three subsets specifically tailored for (i) individual re-identification, (ii) animal pose estimation, and (iii) instance segmentation. Each subset is based on a different set of data; however, they largely overlap. The instance segmentation set includes the majority of data, where each image is paired with a pixel-level mask outlining the lynx’s body, enabling the training of segmentation models. For re-identification, we include only images in which individual coat patterns are clearly visible and where human experts have confirmed identities. The pose estimation subset is slightly smaller due to the labor-intensive annotation process, but also includes challenging samples with unknown identities, such as rare poses, occlusions, and low-visibility conditions. Where appropriate, the image data are accompanied by metadata, such as location, time of observation, viewpoint, and identity, to support model development and evaluation across various ecological scenarios.
Note: Due to the risk of revealing exact animal locations from background details, only the individual re-identification subset with cropped animals is publicly available. The pose estimation and instance segmentation subsets are available upon request111 Access to the dataset requires signing a Non-Disclosure Agreement (NDA) with Friends of the Earth, Czech Republic due to animal safety concerns.
Re-identification: This subset includes 37,400 manually annotated images of 219 individual Eurasian lynxes and 100k synthetic images. The real images were continuously collected for more than 15 years and originate from two primary sources: Hnutí Duha (at Beskydy and Šumava locations) and Šumava National Park (at the Šumava location). Only images with clearly visible flank, leg, or facial patterns (usually used in expert-based manual identification) were selected. The subset includes spatial and temporal metadata to allow the evaluation of identity recognition models under a large domain shift. It also supports evaluation protocols using geo-aware and time-aware splits.
Animal-pose [upon-request]: The pose estimation subset consists of around 15k real images and over 100k synthetically generated images, both annotated with up to 20 keypoints per individual. Due to the scarcity of naturally occurring diverse poses in camera-trap images, the real subset primarily contains side-view walking postures. To counterbalance this, the synthetic subset was generated using a custom Unity-based pipeline that incorporated diverse animations, viewpoints, and textured identities. This enables evaluation of generalization across both pose diversity and realism. The annotations are provided in a standardized COCO-format JSON file. Only images with a single individual are considered.
Instance Segmentation [upon-request]: This subset includes around 40k images annotated with binary instance masks delineating the full body of each lynx. Real images are manually labeled using CVAT, while synthetic images include automatically generated pixel-perfect masks derived from 3D rendering passes. The segmentation masks enable training of detection models capable of distinguishing the lynx from complex, natural backgrounds, including snow, forest undergrowth, and shadows. Each sample includes the associated mask in RLE format, aligned spatially with its corresponding image.
3.2 Predefined Splits
To support robust evaluation under real-world constraints, the CzechLynx dataset is partitioned into three distinct data splits: one geo-aware and two time-aware. Each split is designed to test model generalization across spatial or temporal domains and is provided with a clear training-validation-test split that reflects practical conservation and ecological monitoring scenarios. The summary of sample counts and individual identities per split is presented in Table 6.
Geo-aware split: This split is designed to evaluate spatial generalization. By selecting training data collected from the Hnutí Duha (Beskydy) region and testing on data from the geographically distinct Šumava National Park. The individuals in the training and test sets are disjoint, ensuring that the evaluation reflects out-of-distribution performance across separate ecological contexts. This setting simulates the challenge of deploying models trained in one region to monitor populations in another, which is highly relevant in wider scale (e.g., European) conservation monitoring.
Time-aware open-set split: In this split, data is divided based on time, with the majority of individuals in the test set overlapping with those in the training set. However, test images are drawn from a later time period, thus simulating realistic scenarios where individual appearances might have changed due to seasonal variation, aging, or camera angle. This requires models to generalize to natural visual drift and environmental changes.
Time-aware closed-set split: This split also divides data based on time but ensures that all individuals in the test set are present in the training set. This configuration focuses on evaluating how well models maintain performance over time when all identities are known during training. It reflects practical deployments in ongoing monitoring programs where the same individuals are observed across multiple seasons or years.
# images # identities # locations
training test training test training test
Geo-aware-open 19,751 17,689 150 69 396 347
Time-aware-open 27,959 9,481 197 79 667 292
Time-aware-closed 28,631 8,809 219 213 699 497
Table 6: Number of identities and images in three provided splits; two time-aware and one geo-aware.
4 Technical Validation
To ensure the technical quality of the CzechLynx dataset, we performed a series of validation steps focused on the consistency, accuracy, and usefulness of the data for both computer vision and ecological research.
Annotation Quality Control. The dataset was annotated using standard tools such as CVAT. To reduce labeling errors, a dual-pass validation strategy was used. First, annotations were created or verified by trained annotators. Then, a second round of validation was performed by a separate team to ensure consistency in identity assignment, keypoint placement, and segmentation masks.
Data Consistency and Identity Validation. The re-identification subset includes only images where individual identity could be confirmed by at least two independent observers, based on distinct coat patterns on limbs and flanks. Any ambiguous cases were excluded. This process ensured a high level of reliability in the identity labels across the dataset.
Synthetic Data Comparison. To address pose diversity limitations, a synthetic subset was generated using a 3D lynx model and the Unity engine. We performed side-by-side comparisons of real and synthetic images (see Figure 4) and confirmed that the synthetic data closely mirrors real-world camera trap imagery. The synthetic samples also include complete metadata, such as keypoints, segmentation masks, and identity annotations.
Data Distribution and Coverage. The dataset spans more than 15 years and covers over 219 individual lynx across two populations and multiple regions. Images were collected from over 743 locations, with the majority sampled systematically in a 10$\times$10 km grid. This ensures representative coverage of habitats and individual variation. The dataset also includes time-aware and geographic splits, which allow robust open-set and closed-set evaluation.
Code Availability
The dataset and baseline have been made available through the WildlifeDatasets library. We also include baseline notebooks on the Kaggle in a Code section.
Acknowledgements
This research was supported by the Technology Agency of the Czech Republic, project No. SS05010008. We would like to express our gratitude to Peter Drengubiak, Martina Dušková, Šárka Frýbová, Martin Gendiar, Beňadik Machciník, Michal Králik, Michal Kudlák, Leona Marčáková, Martin Špilák, Zdeněk Tyller, Gabriela Váňová and the dedicated volunteers of Carnivore Tracking Project for their help collecting data and fieldwork.